000 km.) 156 x 10 299 x 1019 2.95 x 1019 2.99 x 10 19 2.96 x 10 19 4.50× 1012 591 x 1012 7.36 x 1022 1.74 x 10
Answer #1

a)

Gravitational force between two objects with mass m1 and m2 separated by distance r is

F=GM1M2/r2

Therefore

F=(6.673*10-11)(7.36*1022)(1.991*1030)/(1.492*1011)2

F=4.4*1020 N

b)

F=(6.673*10-11)(7.36*1022)(5.98*1024)/(3.84*108)2

F=2*1020 N

c)

F=(6.673*10-11)(1.991*1030)(5.98*1024)/(1.496*1011)2

F=3.55*1022 N